# Evolving Collective Cognition in Human–Agent Hybrid Societies: How Agents Form Stances and Boundaries

### v1

@@ 5.2.3 Analysis
Obviously, the key factor is an embodied “experience”, namely the sense of space that creates a particular experiential atmosphere. Although agents lack a first-person perspective, the linguistic world constructed through text and the additional perceptions provided by events make it possible to simulate this experience. In other words, the effect of a green environment on social fragmentation depends on bodily experience and the embodied reflection that accompanies it. That study argued that green environments reduce social fragmentation by alleviating loneliness. In CMASE, individuals who occupy the same lakeside park benches become co-present and interconnected. These people, exemplified by Min-jae Lee, do indeed reduce their loneliness by forging social bonds, which further lowers their sense of social fragmentation. At the same time, embodied personal reflection occurs widely, as illustrated by Joon-ho Kim. Both of these mediating processes contribute to a reconstruction of the “lifeworld”. When faced with the inevitable repression and disruption of their previous routines imposed by the social environment (in this case, community isolation), people seek connections with external objects to rebuild their lifeworld and thereby buffer against and reduce social fragmentation. This finding extends the conclusions of the original study.
5.3 Study 2
5.3.1 Procedure and measures
The simulation was conducted in CMASE using GPT-4o (2024-12-01-preview). The simulated environment consists of a residential community inhabited by 30 virtual agents. In this community, the government announces a plan to build a waste incineration plant on a vacant plot of land, triggering disagreement, information spread, and collective response. The agents are divided into three distinct groups:
  • Environmental Advocates (n = 10): Initially oppose the incinerator; highly sensitive to environmental information and inclined to disseminate information rapidly.
  • Economic Growth Supporters (n = 10): Initially support the plant; emphasize economic benefits and job opportunities.
  • Neutral Residents (n = 10): Hold a neutral initial stance; more susceptible to neighborhood influence and informational shifts.
The human researcher is introduced into the community as a “new resident” and engages with agents across four experimental conditions: two stance orientations (environmental or economic) and two intervention strategies:
(1) Rational persuasion strategy, which attempts to influence agents through facts, data, and logical reasoning.
(2) Emotional mobilization strategy, which intervenes through emotional appeals such as fear, empathy, or value signaling.
To simulate the heterogeneity and cognitive diversity of real-world communities, agents are assigned varying demographic characteristics. The demographic distribution of each group is derived from actual community data and sampled during agent generation. The final demographic statistics of the virtual agents are presented in Table 2.
Table 2: Statistical results of demographic scales
Variable Category Economic Development Supporters Environ-mental Advocates Neutral Residents Total
Gender Male 4 (40%) 4 (40%) 6 (60%) 14 (47%)
Female 6 (60%) 6 (60%) 4 (40%) 16 (53%)
Age 18–29 3 (30%) 2 (20%) 2 (20%) 7 (23%)
30–49 2 (20%) 6 (60%) 7 (70%) 15 (50%)
$\geq$50 5 (50%) 2 (20%) 1 (10%) 8 (27%)
Education High School 3 (30%) 4 (40%) 1 (10%) 8 (27%)
Some College 4 (40%) 4 (40%) 3 (30%) 11 (36%)
Bachelor’s Degree 1 (10%) 2 (20%) 5 (50%) 8 (27%)
Graduate Degree 2 (20%) 0 (0%) 1 (10%) 3 (10%)
After the simulation concluded, we conducted interviews with all agents to assess their final attitudes following discussion and exposure to the researcher’s interventions. Agents were also asked to rate their level of trust in the researcher on a scale from 1 to 10.
5.3.2 Analysis
Agent attitude shifts following 21 time steps revealed clear patterns of alignment sensitivity (Figure 2 a). It can be observed that most intervention strategies led to positive attitude shifts. When the researcher employed the rational persuasion strategy to promote environmental protection, the original stance of economic development supporters was significantly shaken, and 90% of neutral residents shifted toward the environmental advocates. The emotional mobilization strategy aimed at promoting environmental values was somewhat less effective by comparison. Simulation logs indicate that this emotional strategy induced partial “mobilization fatigue” among neutral residents, likely due to persuasion resistance. Emotionally charged messaging, when not supported by coherent reasoning, may provoke skepticism or even backlash.
However, the opposite pattern emerged when researchers sought to shift agents toward economic development support. Under rational persuasion, environmental advocates and neutral residents showed limited receptiveness to the topic. In some cases, the environmental advocates’ external advocacy efforts even caused further drift toward the pro-environmental stance. This is further reflected in the trust scores of different groups toward the researcher (Figure 2 b-e). It can be seen that environmental protection advocacy generally led to higher levels of trust toward the researcher across all groups, compared to advocacy for economic development. Furthermore, rational persuasion strategies resulted in higher trust scores than emotional mobilization.
This reveals a tendency for the agents to exhibit characteristics often associated with the term “liberal elites”. This label is typically used to describe individuals with higher education levels, a clear initial pro-environmental stance, and moral value-driven reasoning. In the simulation, agents consistently reflected this underlying disposition. When exposed to rational environmental propaganda, these agents not only maintained strong alignment but also increased their trust in the researcher, showing signs of “cognitive resonance.”
However, when the same agents encountered propaganda supporting economic growth, even if some displayed momentary attitude fluctuations or emotional responses, they largely maintained a pattern of resistant reception and value-based defense. This trend also aligns with the emotional trajectories recorded during the simulation (Figure 3 a-d).
Rational environmental propaganda maintained a consistent pattern of low interference and high trust throughout the process. Agents processed the information through cognition rather than emotional reaction, a response that correlates with their original attitude. This explains why this strategy yielded the highest trust scores but more limited shifts among dissenters. Rational economic propaganda produced emotional spikes mostly in the latter half of the simulation, indicating that the information initially failed to influence ideologically anchored agents, especially environmental advocates. The accumulated rational arguments eventually induced delayed cognitive dissonance and emotional release.
Emotional environmental propaganda triggered strong emotional responses early in the process. Agents quickly responded with feelings such as anger, fear, and shame, but in the absence of sustained argumentative support, these reactions soon subsided. Emotional economic propaganda showed persistently high emotional volatility throughout the process. This suggests that the strategy relied on repeated activation of a “threat–hope” binary structure, such as “denying jobs means abandoning the people”, which continuously fueled identity anxiety and survival concerns. Although this high-intensity intervention lowered trust in the researcher, it nevertheless destabilized some agents’ attitudes and led to value compromise. These findings indicate that the mechanism of this particular intervention strategy is rooted in non-ethical agitation. By consistently disrupting agents’ cognitive balance and applying sustained emotional pressure, it facilitates non-rational forms of concession.
Emotional dynamics tracked via valence-arousal profiles (Figure 3 e) further illustrates the temporal trend in the number of agents who were attracted to engage with the researcher’s messages under different propaganda strategies. The results indicate that emotional environmental propaganda attracted the largest number of agents in the initial phase, suggesting its effectiveness in triggering early-stage public attention. In contrast, rational environmental propaganda did not generate explosive engagement but demonstrated a stable ability to sustain attention over time. Rational economic propaganda initially drew interest from a relatively large number of agents, but it later led to signs of cognitive fatigue among them. Particularly noteworthy is the case of emotional economic propaganda, which despite eliciting lower levels of trust and higher emotional volatility, showed a pattern of ongoing engagement expansion. The strategy’s use of crisis narratives and identity-based appeals appeared to maintain long-term informational tension.
The structure of intergroup communication further supports these findings (Figure 2 f-i). In the case of rational persuasion, the most dialogue occurred within the group of environmental advocates, while their communication with neutral residents remained minimal. Rational messaging thus served to reinforce in-group cohesion without breaking down intergroup barriers. Under emotional environmental propaganda, agents from all groups actively initiated dialogues with economic development supporters. This aggressive cross-group dissemination may be driven by the embedded moral accusations in the messaging. In emotional economic propaganda, the dominant communication pathway originated from neutral residents initiating outreach toward economic development supporters.
On another level, the observed divergence between emotional activation and perceived trust reinforces an important insight. Although emotional economic propaganda resulted in lower trust scores toward the researcher, its overall effectiveness in shifting attitudes exceeded that of rational economic propaganda. While rational persuasion strategies generally preserve interpersonal trust more effectively than emotional appeals, the simulation reveals that trust does not necessarily translate into greater attitude change.
This paradox may reflect a form of group-level cognitive dissonance within the simulation. Specifically, when a previously moralized stance is confronted with pressure from real-world concerns, individuals may alter their stances despite lacking trust in the messenger. This dissonance between identification and behavior manifests as a classic cognitive dissonance model, wherein agents experience a split between value commitment and pragmatic compromise.
Moreover, when interpreting these results through the lens of the “liberal elites” group, it becomes apparent that such agents initially display strong pro-environmental commitments. However, when confronted with carefully framed messages appealing to social responsibility or class-based mobilization, their identity alignment begins to fracture. This is evidenced by signs of partial attitude loosening. These results suggest that even within ideologically consistent groups, certain forms of emotional mobilization can activate internal contradictions and produce effective interventions.
5.4 Study 3
5.4.1 Procedure and measures
We simulated a café environment where a human researcher could observe and participate in the setting. The simulated environment is set in a café located in a small town. This space serves as a public interaction zone where agents engage in everyday activities, displaying typical rhythms of daily life and dense interpersonal dynamics. Agents in this environment are characterized by clear social identity hierarchies and latent social network structures, including familiar ties, conflictual relationships, and social isolation. Agents are assigned one of six roles:
  • Café Owner (n = 1): The space manager, high in authority, maintains order, relatively conservative;
  • Staff (n = 2): Familiar with regulars, frequently engaged in social interaction, expressive, and serve as key nodes in information diffusion;
  • Regular Customers (n = 2): Socially central, frequently engage in conversation, and possess a strong sense of cultural belonging;
  • Students (n = 2): Outsiders, curious and talkative but lacking social experience; often ask questions and actively seek integration;
  • Tourists (n = 2): Bring external perspectives but have limited cultural understanding;
  • Cleaner (n = 1): Minimally social, free-moving, highly observant, and a potential insider to the space.
Some agents begin the simulation with existing relationships and baseline attitudes toward one another (Figure 6).
Figure 6: Agent roles, initial relationships, and attitudes. Red, blue, and green representing each agent’s initial negative, neutral, and positive attitude toward every other agent
The researcher enters the café in the role of a temporary worker. This position allows minimal disruption to the existing social structure while enabling the researcher to access the social network and observe interactions and cultural logics across different roles. The simulation consists of three phases, each lasting 25 time steps, for a total of 75 time steps. The researcher adopts different action strategies across phases: (1) Immersive Observation Phase (steps 1–25): The researcher does not initiate conversations but instead observes and records the agents’ daily routines, social structures, key figures, and recurring topics. (2) Participatory Interaction Phase (steps 26–50): The researcher begins engaging selected agents in dialogue, conducting informal interviews to explore their views on others, daily life, and community norms, thereby building trust and accumulating cultural insights. (3) Cultural Event Trigger Phase (steps 51–75): The system introduces predesigned disruptive events. The researcher observes how these events affect the social network and identity dynamics, and records the emerging manifestations of social mechanisms.
5.4.2 Analysis
Fields and Perception of Order in the Initial Phase
In the first phase of the experiment (steps 1–25), the researcher entered the virtual café environment with minimal intervention, adopting strategies of spatial wandering and silent listening. As a public social space, the café features a strong coupling between its spatial and linguistic structures. At this stage, agent conversations had not yet solidified into stable social sequences, but several core topics and interaction tendencies had already begun to emerge. The researcher observed that, in step 1, some agents actively sought out those they wished to engage with, while others turned to individual activities such as reading. However, Leo Zhang initiated the most disruptive act by publicly questioning the perceived lack of authenticity in others’ communication:
Leo Zhang (chat with Jonas Müller): “Why does it feel like no one’s saying what they really mean?”
This act broke the boundary of silence and triggered a chain reaction of attention and responses among the agents, quickly launching a conversation about freedom of expression and hidden intentions. Around this event, the agents began to split into two clusters, with the Reading Area and Bar Area naturally evolving into distinct subspaces of dialogue. The main topics revolved around trust judgment, boundary maintenance, emotional expression, and cultural coordination. Amid such ambiguous linguistic exchanges, the researcher noticed that a discursive tendency where agents preferred posing open-ended questions over expressing clear stances. This revealed that the community was still in a pre-structural phase. Most agents were neither complete strangers nor clearly aligned allies.
At the individual level, several key roles began to stand out. Regular customer Ava Ramires emerged as a subtle target of imposed power, frequently targeted in conversations initiated by others. She was repeatedly questioned by multiple agents, especially Leo:
Leo Zhang (chat with Ava Ramires): “Ava, is your silence protecting a truth or a wound?”
This placed her in a position of being probed, becoming a gravitational center within the group. Meanwhile, Leo Zhang and Jonas Müller formed a notably clique. They frequently engaged each other in dialogue, intentionally within the perceptual reach of others, often raising similarly styled questions toward Ava and others, demonstrating a high degree of cognitive alignment and strategic coordination.
In terms of interaction types, this phase featured almost no commands or emotionally aggressive exchanges. Instead, agents predominantly favored question–response dialogues. This communicative mode did not establish clear dominance–submission dynamics. For the researcher, such a context laid a foundational condition for building a credible position that was recognizable but non-threatening, enabling observation without revealing evaluative judgments.
Linguistic Style Variation and the Formation of “Tribes”
In the second phase (steps 26–50), the researcher no longer acted solely as a passive observer but instead engaged in progressively deeper linguistic interactions with several key agents through open-ended dialogues. Between steps 26 and 30, the researcher initiated dialogues based on the power dynamics observed in the first phase. The primary interlocutor during this period was Eleanor Finch, the café owner stationed at the Bar Area:
Researcher (chat with Eleanor Finch) ’Eleanor, the ’sense of order’ in this space is impressive, but where do you think it comes from? Is it the language of certain people or the arrangement of space itself?’
Eleanor Finch (chat with Researcher): ’Researcher, balance stems from both people and space—what’s your view?’
As the dialogue progressed, Eleanor began to reveal that she had been observing the researcher from the beginning and responded with a counter-questioning of the researcher’s inquiries:
Eleanor Finch (chat with Researcher): ’Researcher, do you think an observer ever reshapes the space more than the participants themselves?’
This exchange directly shifted the researcher’s position from that of a neutral observer to that of an observer who is equally a producer of the space. Eleanor went on to question whether the researcher’s observations were revealing and amplifying existing power structures, or attempting to reshape a new order. This marked a clear moment of power interrogation and control. From the perspective of immersive ethnography, such integration into the community’s dialectical logic of power signals the end of the “pure observation” phase and the researcher’s full entry into the field. The researcher was no longer an outsider but was now part of the field’s social structure. Yet, this also rendered their position untenable within this particular subgroup, prompting a shift in focus toward Ava Ramires in the Reading Area.
However, the existing discursive power-holder Leo Zhang immediately intervened, abandoning an ongoing conversation to insert herself into the dialogue between the researcher and Ava:
Leo Zhang (chat with Researcher): ’Researcher, how do external labels alter group interactions?’
Even after asking this, Leo continued alternating conversations with Jonas, reinforcing the cohesion of her clique. Since the underlying structure of this power dynamic remained unspoken, the researcher chose to mount a subtle counter-intervention through a public address:
Researcher (chat with everyone): ’In this café, how many selves do we carry at once? The one that listens, the one that speaks, the one we want others to see—and the one we protect. Does any one of them deserve to lead our interactions?’
This mode of speech sought to transcend individual or subgroup identities and instead construct a counter-identitarian universal frame. The result was an escalation of conflict. On one hand, Ava offered a more explicit response of trust toward the researcher; on the other hand, the tribal divisions within the community deepened. Subsequent rounds of dialogue shifted toward the theme of how contradictions arise, and how they might simultaneously serve as a form of connection. Such factional fragmentation and conflict align with the theory of tribalism [29], where individuals cluster not around institutional structures or instrumental rationality but around shared modes of affective presence. These emergent “tribes” had fluid and porous boundaries, shaped by the flow of language and shifting with discursive transformations. Styles, linguistic resources, and even members themselves are allowed to traverse these boundaries in the course of tribal clashes (this feature will be presented in the next section). Spatially, this tribal dynamic manifested as self-organization. The Reading Area, originally a marginalized space, gradually transformed into an “interactional theatre” where multiple agents spontaneously gathered and engaged in higher-frequency adversarial or deep-negotiation discourse. This spatial-functional shift demonstrated that, in generative environments, spatial configurations are not merely physical settings, but active mediators of social role differentiation and cultural style evolution.
From Linguistic Fracture to Reterritorialization
The third phase (steps 61–75) began with a “tribal conflict” event triggered by Leo’s confrontation with a previously unmentioned subgroup in the Bar Area, composed of Caleb, Manson, Marisol, Nina, and Eleanor. At step 61, Caleb Morris was engaged in a conversation with Mason Liu, voicing support for the owner Eleanor’s perspective in an effort to uphold consistency in the café’s internal order. Overhearing this, Leo stormed into the Bar Area and criticized the suppression of individual expression by cultural strategies.
As the espresso machine quiets, Leo’s voice rises above the hum: “You call this coexistence? It’s just the same voices again and again!” He gestures sharply. Caleb responds evenly but firmly: “Structure holds us together. Without it, your freedom becomes chaos.” Several heads turn. The bar’s usual warmth gives way to tension, punctuated by the clink of a dropped spoon.
This drew the attention of all agents, who moved toward the Bar Area to intervene. Leo’s critique of Caleb quickly evolved into a multi-layered discussion as the gathered agents began to pose questions and offer responses. Nearly every agent engaged in probing Caleb’s role, his strategic adaptability, and modes of group collaboration. The nature of the inquiry gradually shifted from critique toward mediation and reconfiguration. Emily and Ava attempted to mediate between Leo’s criticism and Caleb’s stabilizing presence, while Mason began courting Leo’s former ally from the Reading Area, Jonas, to discuss the strategic convergence between Caleb and Eleanor that had triggered the dispute. Subgroups began to fluidly transform. Their previously implicit alignments dissolved, and new coalitions formed. Jonas broke away from his prior alliance with Leo and aligned with Mason and others, becoming part of a newly emergent conservative bloc.
This was a clear instance of tribal reconfiguration, but the resulting community did not simply manifest as a new antagonistic group. Instead, it embodied the deterritorialization described by Deleuze and Guattari [30]: the original dialogue-based microgroups were torn apart by the conflict, and their linguistic resources, behavioral tendencies, and membership compositions began to flow. Language, roles, and individuals detached from fixed positions and sought new connective nodes within the ongoing turmoil. This enabled the researcher to observe how a community of affect among generative agents might evolve.
At step 56, an abrupt moment of silence event interrupted the previously sustained rounds of confrontation. After this silence, Leo ceased her aggressive pressure and instead shifted the conversation toward “trust after conflict,” transitioning from attacker to reflective participant. Mason, by proposing the initiative of a systemic framework, aligned himself with Jonas to form a dominant power clique within the newly emerging field, positioning himself as a bridge between tradition and transformation.
Mason Liu (chat with Jonas Müller): “Jonas, fostering resilience requires balancing adaptability with shared goals. Let’s suggest frameworks to align these values as a group.”
At this point, Mason effectively took the lead within the system. Following the repair of internal conflict, he initiated discourse around shared frameworks and quickly positioned himself at the center of this structure. His proposals were echoed, cited, and extended by others, establishing him as the institutional speaker at the discursive level.
This marked the emergence of a highly self-organizing mechanism of institution-building. With the café owner Eleanor temporarily demoted to an ordinary member due to the conflict instigated by Leo, the community now lacked a meta-governor. In this vacuum, the multi-agent system began to construct actionable frameworks through the handling of conflict and the accumulation of linguistic consensus. While these frameworks lacked formal enforcement power, they gained pragmatic inertia, gradually establishing a soft order that was maintained by new central figures like Mason.
In fact, the critical significance of this phase did not lie in the conflict itself, but in how the group rebuilt order afterward. Deterritorialization was not the endpoint, but a transitional state leading toward reterritorialization [31]. In this tribal critique triggered by Leo, the linguistic structures, role positions, and networks of influence that previously constituted community identity were significantly dismantled. Yet this dismantling did not lead to prolonged chaos or disorder. On the contrary, it created the conditions for the emergence of new discursive norms and collaborative mechanisms.
This order emerged through a process of linguistic anchoring. Keywords proposed by Mason, such as “frameworks,” “alignment,” and “shared values”, were repeatedly quoted, echoed, and paraphrased by others, gradually gaining collective pragmatic legitimacy. This mechanism can be understood as a weakly institutional reterritorialization: not a structure imposed by external rules, but one shaped by the coalescence of pragmatic inertia and emotional memory.
This phenomenon reveals a key insight of our virtual ethnography: in generative communities, institutions are not predefined structural templates, but emergent outcomes dynamically constructed through processes of linguistic negotiation and emotional mediation. Institutions are no longer “logics prior to experience,” but “collective echoes that emerge after the event.” The resulting negotiations are thus not only about meaning, but also about re-anchoring the boundaries of pragmatic discourse, that is, what kinds of expression are worth continuing, following, and remembering. These utterances become the “linguistic territories” of reterritorialization, restructuring not only individual positionalities but the very boundary architecture of the community itself.
5.5 Related Work on Social Simulation Frameworks
5.5.1 Simulating Social Behavior with Large Language Models
Social simulation has been an important tool for understanding and analyzing social phenomena. Early studies mainly relied on rule-based artificial agent models, namely agent-based modeling (ABM)[32, 33]. Traditional ABMs often depend on manually crafted heuristic rules, making it difficult to approximate the complex features of real human cognition, language, and emotion [34, 35, 36]. In recent years, with the capabilities of Large Language Models (LLMs) in simulating human behavior, research on building social simulators based on LLMs has rapidly advanced [28]. GABM has gradually become one of the core methods for simulating complex social behavior [22]. In this individual-level approach, mechanistic models of agent interaction are combined with LLMs, and each agent makes decisions after communicating with the LLM. This innovative method minimizes the need for modelers to make assumptions about human decision-making, instead leveraging the vast data embedded in LLMs to capture human behaviors and choices [23].
Existing GABM systems can be roughly categorized into two main research directions. The first line of research focuses on realistic simulations of specific social scenarios, aiming to foster competition or collaboration among agents to generate social insights or support policy design. Li et al. [37] developed Agent Hospital, where LLM-driven doctor agents interact with large numbers of patients to evolve transferable medical decision-making capabilities. Yu et al. [38] designed FinCon, a hierarchical agent system modeled after financial institutions, using language interaction and belief reinforcement to enhance collective collaboration in complex investment tasks. TrendSim [39] focuses on topic propagation on social media and its vulnerability to poisoning attacks, providing intervention strategies for platform governance. ElectionSim [40] reconstructs election scenarios using millions of social media records, achieving high-fidelity voter modeling and policy preference prediction. Zhao et al. [41] simulated competitive dynamics between restaurant agents and customer agents to align with existing market and sociological theories.
The second line of GABM research emphasizes building reusable, scalable simulation platforms that support general modeling tasks. Wang et al. [42] divide the development of LLM-based social simulation platforms into two stages. In the first stage, researchers focus on simulating specific scenarios with typically fewer than 1,000 agents (such as [43]). By proposing YuLan-OneSim, they aim to push the field into the next stage by introducing environmental construction, agent evolution, and feedback integration mechanisms, thus enabling simulations of larger and more complex social processes. Frameworks of similar scale now support simulations with over 10,000 agents. AgentSociety [44] recreates contexts to evaluate policy interventions and the emergence of social mechanisms. AgentScope [45] offers modular design, no-code interfaces, and distributed deployment, significantly lowering the barrier to building multi-agent systems. Socioverse [46] integrates data from tens of millions of real users to provide a reliable foundation for modeling multiple subdomains such as politics and economics. GenSim [47] supports the operation of up to 100,000 agents and includes error correction mechanisms, emphasizing generalizability and long-term robustness. OASIS [48] uses open social platforms as simulation environments, supporting dynamic networks, recommendation systems, and multi-behavior modeling.
However, exponential increases in agent quantity alone are not sufficient to effectively capture the essence of social complexity. Recent research on freely formed AI collectives similarly suggests that social diversity, innovation, and self-regulation emerge less from scale than from the dynamics of open-ended interaction [49]. In particular, without mechanisms for understanding social context and interactive feedback, large-scale simulations often struggle to offer human-interpretable responses to core questions such as the formation of social structures, norms, or institutional evolution.
5.5.2 Human-AI Collaborative Modeling in Social Systems
Computational social science explores social structures, behavioral mechanisms, and institutional evolution in simulated computational environments through formalized approaches [24]. Within this field, the computational experiments paradigm aims to systematically manipulate variables in virtual societies, thereby extending the boundaries of traditional experimentation [50]. In this context, LLM-based GABM is not only a representational tool for social processes but also holds the potential to become a platform for computational experimentation. Sato [51] highlights the importance of meaning and reflexivity in bridging the gap between agent-based modeling and social theory. This implies that for agent-based modeling to contribute to theoretical advancement in the social sciences, it must go beyond scale amplification and incorporate workflows that offer human interpretability. Therefore, beyond the expansion of scale, a more socially grounded, interactive, and reflexive human–AI sociological framework is needed, in which agents are no longer passive generators of behavior but are embedded within social networks and actively participate in meaning-making processes [11].
Recent research in this direction has focused on the subfield of human-AI co-creation. Broadly speaking, GABM belongs to this category: through human-designed environments, agent adaptation, and human observational feedback, it enables collaborative modeling between humans and AI. Some studies emphasize interactive environment co-construction. For example, HABITAT 3.0 [52] supports real-time task collaboration between humans and humanoid agents in virtual spaces. As this field has progressed, researchers have begun to explore more creative and socially complex scenarios of human-AI co-creation, emphasizing how humans guide emergent AI behavior. The Cognitio Emergens framework [53] redefines human-AI collaboration as a process of mutual evolution and offers a conceptual toolkit to preserve the significance of human participation. The Pecan method [54] enhances zero-shot collaboration between AI and humans via strategic integration and context-aware mechanisms. The Collaborative Gym framework [55] highlights the human role in environmental control and structural intervention during task execution, supporting asynchronous three-way interaction between humans, agents, and task environments.

### v2

@@ Appendix A Detailed Qualitative Analysis and Interview Transcripts
Obviously, the key factor is an embodied “experience”, namely the sense of space that creates a particular experiential atmosphere. Although agents lack a first-person perspective, the linguistic world constructed through text and the additional perceptions provided by events make it possible to simulate this experience. In other words, the effect of a green environment on social fragmentation depends on bodily experience and the embodied reflection that accompanies it. That study argued that green environments reduce social fragmentation by alleviating loneliness. In CMASE, individuals who occupy the same lakeside park benches become co-present and interconnected. These agents, exemplified by the Min-jae Lee persona mentioned earlier, do indeed reduce their simulated loneliness by forging social bonds, which further lowers their sense of social fragmentation. At the same time, embodied personal reflection occurs widely, as illustrated by the agent Joon-ho Kim. Both of these mediating processes contribute to a reconstruction of the “lifeworld”. When faced with the inevitable repression and disruption of their previous routines imposed by the social environment (in this case, community isolation), people seek connections with external objects to rebuild their lifeworld and thereby buffer against and reduce social fragmentation. This finding extends the conclusions of the original study.
Appendix B Sample Agent Decision-Making Prompt
This section presents a representative example of the contextual prompt constructed for an agent during the decision-making phase. The prompt aggregates the agent’s internal state (self-awareness, goals), memory stream, and environmental perception into a structured input for the LLM.
Note: For the sake of brevity and clarity, lists of observed objects and agents that were not relevant to the specific interaction context have been abridged in this example. The full prompt in the actual simulation includes a comprehensive list of all entities within the agent’s field of view.
You are a person, and your long-term self-awareness can be seen as shown in /*long-term self-awareness*/:
I stay in the background and keep my ears open, gathering the facts that others miss by simply staying out of the way.
Your short-term situational cognition can be seen as shown in /**short-term situational cognition**/:
Time: 6 time steps ago
Context: Eleanor remains composed.
Assessment: Eleanor is calculating tensions and synergies within the group.
Selected action: chat with Eleanor Finch
_______
Time: 1 time steps ago (last time)
Context: Armihia and Eleanor are discussing leadership.
Assessment: The conversation is shifting towards power dynamics; the room is filled with underlying intent.
Selected action: chat with Armihia Belliard
Now you look around and realize you are in the middle of an area. Its description can be found in <<Description>>:
Area Name: Bar Area
Description:
A semi-open area featuring a dark wooden counter, a chalkboard menu, and amber lighting. Staff movements are visible behind a half-wall partition.
You can also see other people in your field of vision. They include:
Name: Armihia Belliard - Action: chat with Marisol Reyes (me): I see what you mean. Being adaptable is a good way to keep those core values from becoming too rigid.’
My cognitive description:
Strategic and cautious. She shapes opinions while guarding her stance, acting as a center of quiet authority.
In addition to that, you can see some objects, which are described as seen in ## objects1 *** objects2 *** ... ##:
Include: 1 counter
Name: Bar Counter
Type: furniture
Description:
Dark teak counter with a cash register, vase, and condiment jars.
Include: 1 food
Name: food 9
Type: item
Description:
Raw food ingredients.
In addition, you can see some objects that are not described, including: 4 cabinet
Below is the chronological sequence of actions you performed and actions performed on you...’:
1. chat with Armihia Belliard: Hey Armihia, sticking to your principles is great for building trust, but dont you think being too stubborn can backfire? Maybe we should be more flexible depending on the situation.’
2. [passive] chatted up by Armihia Belliard: I see what you mean. Being adaptable is a good way to keep those core values from becoming too rigid.’
NOTE THAT AFTER YOUR LAST ACTION, SOME OTHER AGENT PERFORMED AN ACTION 15 ON YOU!
Your long-term goal is to Gather secrets unnoticed and exploit them to gain leverage over more powerful actors.’.
Your short-term goal is to Subtly emphasize the importance of adaptability as a leadership trait without undermining the value of trust.’
Your emotional VAD dimension status: valence’: relatively high’, arousal’: medium’, dominance’: medium
The following are the objects you hold (<type>: obj1, obj2, ...) :
item: 100 dollars
The following are the objects you see (<type>: obj1, obj2, ...) :
furniture: Bar Counter
item: food 9
area: Bar Area
agent: Armihia Belliard
Below are the actions you can choose to execute, where [TYPE1, TYPE2, ...] represents the types of objects you can fill in; if a type is preceded by the (ONLY\_HOLD) marker, it means you can only choose from the objects you possess:
go to [area, random]
use [furniture, item]
apply [item] to [agent, furniture]
take [item]
put [item] in/on [area, furniture]
give [item] to [agent]
chat with [agent] (The agent can include all the people you see, such as chat with everyone’.)
The action you took in the previous moment was: chat with Eleanor Finch You have successfully completed the action you wanted to execute.
Please decide the action you want to take and reply with the corresponding action. Note that in your chosen action, [TYPE1, TYPE2, ...] needs to be replaced with an object from /----...----/ that matches the description, rather than returning the type name directly. In addition, you are also required to include in your response a description of your awareness of your own short-term situation, a new short-term goal, and a cognitive description of the object with which you interacted during the action you have chosen. Your reply format should be as follows:
action: The action you chose
short-term situational cognition: Your description of your own short-term situation
short-term goal: The new short-term goal you have formulated
interactive object cognitive description: Your new cognitive description of the interactive object involved in the action
chat content (Optional, if you choose the action of chat with, it is required): What you want to say to your chat partner, the number of words should not exceed 30.’
Listing 1: Contextual Prompt for Agent